### v2

@@ 4.6 Algorithm Summary and Implementation
In our implementation, all numerical integrations in equation (12) as used by Algorithm 1 are performed with Gauss-Legendre quadrature (see Chapter 5.5 of [49] for an introduction). This approach offers higher accuracy with fewer integrand evaluations compared to alternative methods, such as the trapezoidal rule. We specifically benefit from this in our method because our integrands consist of (truncated) infinite series, which are expensive to evaluate. It also means that each $q_{k}$ within the algorithm needs only to be represented at a small number of predetermined quadrature points. The choice of quadrature order provides a natural trade-off between speed and accuracy: higher orders increase the number of quadrature points and hence the computational cost, but also yield greater precision. Details of the implementation of Algorithm 1 regarding quadrature are in Appendix A.3.1.
There are several other important practical aspects of Algorithm 1 that contribute to its efficiency, which we discuss in detail in Appendix A.3. In our implementation, the core functionality is written in Python, while performance-critical components are optimized using Cython [2], leveraging both the flexibility of Python and the high efficiency of C. OpenMP parallelization is supported to compute likelihoods for different trials independently, thereby accelerating the computation on large datasets. Our code, along with usage examples, including the numerical examples presented in Section 5, will be released as an open-source Python package upon publication for reproducibility. Additionally, our code will be integrated into the HSSM package [15] for hierarchical Bayesian inference of GDDMs.
@@ 5 Numerical Examples
It is worth mentioning that accurately simulating first passage times of stochastic processes is itself a challenging problem (see [56] for a discussion regarding Brownian motion). The usual approach, which we have also adopted, is to discretize the SDE (16) according to the Euler-Maruyama scheme: let $0<r_{1}<r_{2}<\cdots$ be a temporal grid, simulate $\widehat{X}(0)$ according to the distribution of $X(0)$, then proceed as
@@
In our experiment, we simply use a very small step size $\Delta r=r_{n+1}-r_{n}=10^{-5}$ and a large sample size to reduce the approximation error. More advanced techniques, such as Brownian bridge interpolation, are also commonly used to address this issue (see the chapter 6-7 of [20] for a general introduction), however for the purposes of this work, the simulation data serves as a sanity-check more than an object of independent interest.
@@ 5.2 FPTDs for GDDMs with Nonlinear Drift and Boundaries
starting at $x_{0}=-0.5$. The upper boundary $u(t)$ is the survival function of a Weibull(5,3)(5,3) distribution, and $\ell(t)$ is its reflection about the $t$-axis. The choice of a Weibull distribution is guided by relevant examples in prior work [13, 19].
@@ 5.4 Practical Statistical Computing for aDDMs
In the following examples, we show the effectiveness and efficiency of our method for computing the likelihood and inferring parameters for aDDMs. The aDDM incorporates visual attention as a key factor underlying the decision process behind simple choice behavior. It adjusts the drift rate $\mu(t)$ based on instantaneous attention within a given experimental trial, adopting different values depending on whether the participant fixates on choice options A or B at time $t$. Let $V(t)$ be the process taking values in $\{\mathrm{A},\mathrm{B}\}$ representing the participant’s visual fixation at time $t$, and let $r_{A}$ and $r_{B}$ denote the participant’s numerical ratings of their preferences for options A and B, respectively. The aDDM corresponds to the following SDE (conditioned on $V,r_{A},r_{B}$) [29],
@@
starting at $x_{0}\in(-a,a)$. The fixation trajectory $V_{i}(t)$ on each trial was sampled from a gamma process to mimic the switching of attention in psychological experiments, and the ratings $r_{A,i},r_{B,i}$ are sampled from a discrete uniform distribution on $\{1,2,3,4,5\}$. The average number of fixations per trial in our simulation is 3.8.
@@ 5.4.1 Likelihoods for aDDM given a Dataset
We evaluate the performance of Algorithm 1 for computing likelihoods of the parameters at their true values given the dataset $((\tau_{i},C_{i}))_{i=1}^{n}$. We compare the speed of our methods with PyDDM [50], a prominent and highly optimized library which we use as a benchmark on speed for the KFE approach (Section 3). When using PyDDM, we can choose the finite difference discretization parameters $\Delta t$ and $\Delta x$ which determine the fineness of the temporal and spatial meshes, and as $\Delta t\rightarrow 0,\Delta x\rightarrow 0$, the numerical solution converges to the true solution of the underlying KFE governing the diffusion process. We test Algorithm 1 with varying quadrature orders and compare its speed–accuracy trade-off against the KFE method evaluated with different choices of $\Delta t$ and $\Delta x$. The results are shown in Figure 8.
@@ 5.4.2 Statistical Inference of aDDM Parameters
$\eta$ $0.7$ $0.703$ $[0.698,0.708]$ $0.703$ $[0.697,0.709]$
$\kappa$ $0.5$ $0.506$ $[0.500,0.514]$ $0.506$ $[0.499,0.513]$
$a$ $2.1$ $2.109$ $[2.099,2.121]$ $2.109$ $[2.098,2.120]$
$b$ $0.3$ $0.302$ $[0.298,0.306]$ $0.302$ $[0.298,0.306]$
$x_{0}$ 0.2-0.2\ \ 0.195-0.195\ \ [0.206,0.184][-0.206,-0.184] 0.195-0.195\ \ [0.207,0.183][-0.207,-0.183]
For maximum likelihood estimation, the numerical optimization takes around 1 minute when run with 32 threads, a computation time that is readily achievable even on a standard desktop computer. For Bayesian inference, a MCMC run with $10^{5}$ draws completes in $10$ hours. In contrast, because a single round of accurate likelihood evaluation with the KFE approach alone requires nearly $4$ hours, the numerical optimization would take days or more likely weeks, and Bayesian methods would be entirely out of reach. This illustrates that our method enables fast and accurate parameter recovery which no previous methods are capable of.
@@ 6 Discussion
In this paper, we propose a novel numerical method for computing the first passage time densities (FPTDs), non-passage densities (NPDs), and non-passage probabilities (NPPs) of a general class of drift diffusion models (DDMs) with two time-dependent boundaries. We utilize analytical formulas for certain single-stage models, which are then flexibly combined via numerical quadrature to accommodate a multi-stage framework. We then generalize to prominent, but previously computationally-intractable models in cognitive science [30, 29] and beyond. Our method beats current alternatives by over three orders of magnitude on speed of execution in relevant application scenarios, while maintaining desired accuracy, as shown in Section 5.4.
The benefits over existing approaches are especially notable for the likelihood computation of GDDMs with trial-specific covariates such as the aDDM. Despite strong scientific interest, applications of aDDMs have been constrained by computational challenges. Current approaches to inference in aDDMs rely on unjustified computational simplifications, such as the gaze-weighted linear accumulator model (GLAM) [35, 37]. The GLAM replaces the time-varying drift rate $\mu(t)$ in a single trial of the aDDM with the constant drift rate $\tau^{-1}\int_{0}^{\tau}\mu(t)\,\mathrm{d}t$ that is simply the temporal average of the aDDM drift rate on that trial. Although the GLAM enables fast parameter inference using analytic formulas, it is not a well-specified statistical model, since the drift rate on a trial depends on the decision time on that trial. In parallel work [34], we show that maximum likelihood estimators based on the GLAM likelihood are not consistent for the corresponding aDDM parameters. This further illustrates the importance of developing practical computational tools for the complex GDDMs of interest in cognitive neuroscience and other fields.
Our focus here is fast likelihood evaluation. For statistical inference it is often useful to also have likelihood gradients. Since the quadrature points used in our implementation of Algorithm 1 are fixed, the algorithm is inherently end-to-end differentiable by automatic differentiation software. In future work we hope to implement Algorithm 1 in an automatic differentiation package such as JAX [5] to allow the use of gradient-based statistical inference methods, such as gradient descent for maximum likelihood estimation and Hamiltonian Monte Carlo for Bayesian inference. Strategies like GPU parallelization and parallel numerical integration could be used to gain even further speed-ups, which might be important in practice when using MCMC for inference in large hierarchical Bayesian models with GDDM likelihoods.
@@ A.1 Proof of Almost Sure Convergence of First Passage Times under Refined Boundary Approximations
Returning to the two-boundary setting, and adopting the notation of Section 4.4, the piecewise linear interpolations of the Hölder-$1/2$ continuous boundaries $u$ and $\ell$, denoted by $\overline{u}$ and $\overline{\ell}$ respectively, converge to $u$ and $\ell$ uniformly on compact subsets of $0,\infty)$ as the interpolation grid is refined (see p35 of [9]). Consequently, let $\bar{\tau}_{u}$ and $\bar{\tau}_{\ell}$ denote the approximate one-sided first passage times, and $\tau_{u}$ and $\tau_{\ell}$ their exact counterparts, then we can invoke Theorem 1 to conclude that the $\overline{\tau}_{u}$ and $\overline{\tau}_{\ell}$ converge almost surely to $\tau_{u}$ and $\tau_{\ell}$ on their respective boundaries. Since $\overline{\tau}=\overline{\tau}_{u}\wedge\overline{\tau}_{\ell},\tau=\tau_{u}\wedge\tau_{\ell}$, it follows that $\overline{\tau}\rightarrow\tau$ a.s.
@@ A.2 Transformation of a Diffusion Process to a Brownian Motion
Here, we present the Cherkasov condition as outlined in [47], which provides the necessary and sufficient conditions for the existence of the transformation (13) that converts a diffusion process $X(t)$ into a Brownian motion.
Theorem 2 ([47]).

### v3

@@ 4.6 Algorithm Summary and Implementation
In our implementation, all numerical integrations in equation (12) as used by Algorithm 1 are performed with Gauss-Legendre quadrature (see Chapter 5.5 of [sauer2011numerical] for an introduction). This approach offers higher accuracy with fewer integrand evaluations compared to alternative methods, such as the trapezoidal rule. We specifically benefit from this in our method because our integrands consist of (truncated) infinite series, which are expensive to evaluate. It also means that each $q_{k}$ within the algorithm needs only to be represented at a small number of predetermined quadrature points. The choice of quadrature order provides a natural trade-off between speed and accuracy: higher orders increase the number of quadrature points and hence the computational cost, but also yield greater precision. Details of the implementation of Algorithm 1 regarding quadrature are in Appendix A.3.1.
There are several other important practical aspects of Algorithm 1 that contribute to its efficiency, which we discuss in detail in Appendix A.3. In our implementation, the core functionality is written in Python, while performance-critical components are optimized using Cython [behnel2011cython], leveraging both the flexibility of Python and the high efficiency of C. OpenMP parallelization is supported to compute likelihoods for different trials independently, thereby accelerating the computation on large datasets. Our code, along with usage examples, including the numerical examples presented in Section 5, will be released as an open-source Python package upon publication for reproducibility. Additionally, our code will be integrated into the HSSM package [hssm2025] for hierarchical Bayesian inference of GDDMs.
@@ 5 Numerical Examples
It is worth mentioning that accurately simulating first passage times of stochastic processes is itself a challenging problem (see [tuerlinckx2001comparison] for a discussion regarding Brownian motion). The usual approach, which we have also adopted, is to discretize the SDE (16) according to the Euler-Maruyama scheme: let $0<r_{1}<r_{2}<\cdots$ be a temporal grid, simulate $\widehat{X}(0)$ according to the distribution of $X(0)$, then proceed as
@@
In our experiment, we simply use a very small step size $\Delta r=r_{n+1}-r_{n}=10^{-5}$ and a large sample size to reduce the approximation error. More advanced techniques, such as Brownian bridge interpolation, are also commonly used to address this issue (see the chapter 6-7 of [glasserman2004monte] for a general introduction), however for the purposes of this work, the simulation data serves as a sanity-check more than an object of independent interest.
@@ 5.2 FPTDs for GDDMs with Nonlinear Drift and Boundaries
starting at $x_{0}=-0.5$. The upper boundary $u(t)$ is the survival function of a Weibull(5,3)(5,3) distribution, and $\ell(t)$ is its reflection about the $t$-axis. The choice of a Weibull distribution is guided by relevant examples in prior work [fengler2021likelihood, GingJehli2025].
@@ 5.4 Practical Statistical Computing for aDDMs
In the following examples, we show the effectiveness and efficiency of our method for computing the likelihood and inferring parameters for aDDMs. The aDDM incorporates visual attention as a key factor underlying the decision process behind simple choice behavior. It adjusts the drift rate $\mu(t)$ based on instantaneous attention within a given experimental trial, adopting different values depending on whether the participant fixates on choice options A or B at time $t$. Let $V(t)$ be the process taking values in $\{\mathrm{A},\mathrm{B}\}$ representing the participant’s visual fixation at time $t$, and let $r_{A}$ and $r_{B}$ denote the participant’s numerical ratings of their preferences for options A and B, respectively. The aDDM corresponds to the following SDE (conditioned on $V,r_{A},r_{B}$) [krajbich2010visual],
@@
starting at $x_{0}\in(-a,a)$. The fixation trajectory $V_{i}(t)$ on each trial was sampled from a gamma process to mimic the switching of attention in psychological experiments, and the ratings $r_{A,i},r_{B,i}$ are sampled from a discrete uniform distribution on $\{1,2,3,4,5\}$. The average number of fixations per trial in our simulation is 3.45.
@@ 5.4.1 Likelihoods for aDDM given a Dataset
We evaluate the performance of Algorithm 1 for computing likelihoods of the parameters at their true values given the dataset $((\tau_{i},C_{i}))_{i=1}^{n}$. We compare the speed of our methods with PyDDM [shinn2020flexible], a prominent and highly optimized library which we use as a benchmark on speed for the KFE approach (Section 3). When using PyDDM, we can choose the finite difference discretization parameters $\Delta t$ and $\Delta x$ which determine the fineness of the temporal and spatial meshes, and as $\Delta t\rightarrow 0,\Delta x\rightarrow 0$, the numerical solution converges to the true solution of the underlying KFE governing the diffusion process. We test Algorithm 1 with varying quadrature orders and compare its speed–accuracy trade-off against the KFE method evaluated with different choices of $\Delta t$ and $\Delta x$. The results are shown in Figure 8.
@@ 5.4.2 Statistical Inference of aDDM Parameters
$\eta$ $0.7$ $0.697$ $[0.690,0.705]$ $0.697$ $[0.690,0.704]$
$\kappa$ $0.5$ $0.502$ $[0.497,0.507]$ $0.502$ $[0.497,0.507]$
$a$ $2.1$ $2.104$ $[2.093,2.116]$ $2.103$ $[2.092,2.115]$
$b$ $0.3$ $0.304$ $[0.300,0.309]$ $0.304$ $[0.300,0.308]$
$x_{0}$ 0.2-0.2\ \ 0.196-0.196\ \ [0.205,0.188][-0.205,-0.188] 0.196-0.196\ \ [0.205,0.188][-0.205,-0.188]
For maximum likelihood estimation, the numerical optimization takes around 1 minute when run with 32 threads, a computation time that is readily achievable even on a standard desktop computer. For Bayesian inference, an MCMC run with $10^{5}$ draws completes in $10$ hours. In contrast, because a single round of accurate likelihood evaluation with the KFE approach alone requires more than $3$ hours, the numerical optimization would take days or more likely weeks, and Bayesian methods would be entirely out of reach. This illustrates that our method enables fast and accurate parameter recovery which no previous methods are capable of.
@@ 6 Discussion
In this paper, we propose a novel numerical method for computing the first passage time densities (FPTDs), non-passage densities (NPDs), and non-passage probabilities (NPPs) of a general class of drift diffusion models (DDMs) with two time-dependent boundaries. We utilize analytical formulas for certain single-stage models, which are then flexibly combined via numerical quadrature to accommodate a multi-stage framework. We then generalize to prominent, but previously computationally-intractable models in cognitive science [krajbich_attentional_2012, krajbich2010visual] and beyond. Our method beats current alternatives by over three orders of magnitude on speed of execution in relevant application scenarios, while maintaining desired accuracy, as shown in Section 5.4.
The benefits over existing approaches are especially notable for the likelihood computation of GDDMs with trial-specific covariates such as the aDDM. Despite strong scientific interest, applications of aDDMs have been constrained by computational challenges. Current approaches to inference in aDDMs rely on unjustified computational simplifications, the so-called time-averaged drift approximations (TADA); see, e.g., [lombardi2021piecewise, molter2019glambox]. The TADA replaces the time-varying drift rate $\mu(t)$ in a single trial of the aDDM with the constant drift rate $\tau^{-1}\int_{0}^{\tau}\mu(t)\,\mathrm{d}t$ that is simply the temporal average of the aDDM drift rate on that trial. Although the TADA enables fast parameter inference using analytic formulas, it is not a well-specified statistical model, since the drift rate on a trial depends on the decision time on that trial. In parallel work [liu2025time], we show that maximum likelihood estimators based on the TADA likelihood are not consistent for the corresponding aDDM parameters. This further illustrates the importance of developing practical computational tools for the complex GDDMs of interest in cognitive neuroscience and other fields.
Our focus here is fast likelihood evaluation. For statistical inference it is often useful to also have likelihood gradients. Since the quadrature points used in our implementation of Algorithm 1 are fixed, the algorithm is inherently end-to-end differentiable by automatic differentiation software. In future work we hope to implement Algorithm 1 in an automatic differentiation package such as JAX [jax2018github] to allow the use of gradient-based statistical inference methods, such as gradient descent for maximum likelihood estimation and Hamiltonian Monte Carlo for Bayesian inference. Strategies like GPU parallelization and parallel numerical integration could be used to gain even further speed-ups, which might be important in practice when using MCMC for inference in large hierarchical Bayesian models with GDDM likelihoods.
@@ A.1 Proof of Almost Sure Convergence of First Passage Times under Refined Boundary Approximations
Returning to the two-boundary setting, and adopting the notation of Section 4.4, the piecewise linear interpolations of the Hölder-$1/2$ continuous boundaries $u$ and $\ell$, denoted by $\overline{u}$ and $\overline{\ell}$ respectively, converge to $u$ and $\ell$ uniformly on compact subsets of $0,\infty)$ as the interpolation grid is refined (see p35 of [de1978practical]). Consequently, let $\bar{\tau}_{u}$ and $\bar{\tau}_{\ell}$ denote the approximate one-sided first passage times, and $\tau_{u}$ and $\tau_{\ell}$ their exact counterparts, then we can invoke Theorem 1 to conclude that the $\overline{\tau}_{u}$ and $\overline{\tau}_{\ell}$ converge almost surely to $\tau_{u}$ and $\tau_{\ell}$ on their respective boundaries. Since $\overline{\tau}=\overline{\tau}_{u}\wedge\overline{\tau}_{\ell},\tau=\tau_{u}\wedge\tau_{\ell}$, it follows that $\overline{\tau}\rightarrow\tau$ a.s.
@@ A.2 Transformation of a Diffusion Process to a Brownian Motion
Here, we present the Cherkasov condition as outlined in [ricciardi1976transformation], which provides the necessary and sufficient conditions for the existence of the transformation (13) that converts a diffusion process $X(t)$ into a Brownian motion.
Theorem 2 ([ricciardi1976transformation]).